OpenStudy (anonymous):

Help!! Simplify: Questions below A. 16 B. 6 C. 2 D. 8

OpenStudy (anonymous):



OpenStudy (anonymous):

Step by step or just the answer?

OpenStudy (anonymous):

both please lol :)

OpenStudy (anonymous):

Okay one second lol

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

So basically this is a "cube root" or well that's what I call it It means that you need to find a number that when it is CUBED (so multiplied by itself three times) it will give the number in the square root symbol (so in your case 216) That was a horrible explanation so I'll use numbers What number multiplied 3 times itself makes 216? Let's try 10: so: 10 * 10 * 10 = 1000 So we know the number is LESS than 10 Let's try 5: 5 * 5 * 5 = 125 So we know that the number cubed must be IN BETWEEN 5 and 10 Let's try 6: 6 * 6 *6 = 216 There is your answer! So your answer is B!

OpenStudy (anonymous):

wow thats so much easier then what my teacher told me:p thank you !!!!

OpenStudy (anonymous):

lol no problem :)